OpenStudy (anonymous):

-8 is less than y, and y is less than or equal to 9

jimthompson5910 (jim_thompson5910):

"-8 is less than y" turns into -8 < y "y is less than or equal to 9" turns into \(\Large y \le 9\) So put together, we have \(\Large -8 < y \ \text{ and } \ y \le 9\) which combine to this compound inequality \(\Large -8 < y \le 9\)